# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v2

@@ 6 Limitations
We assume the operational oracle $Y$ aligns with stakeholder values; oracle selection itself is a governance question outside the scope of this work (Assumption A0 in App. C.14).
@@
IPS requires support overlap between $\pi_{0}$ and each $\pi^{\prime}$; DR benefits from overlap but remains consistent via its outcome model when overlap is poor; Direct requires no overlap (it generates fresh responses under $\pi^{\prime}$). When overlap is poor, raw ratios are heavy-tailed and uncertainty inflates. Mitigations: SIMCal–W reduces dispersion and raises $\mathrm{ESS}$; if tails persist we (i) gate on $\mathrm{ESS}$ and Hill indices, (ii) use overlap weighting or cohort restriction, and (iii) run an online check when $\widehat{\alpha}_{\text{Hill}}{<}1$ or single–row dominance persists (App. D).
@@
AutoCal–R assumes mean sufficiency and monotonicity in $S$ (or a learned index). If strained, the two–stage fallback preserves mean honesty but targets $\mathbb{E}[f(S^{\pi^{\prime}})]$. Mitigations: surface reliability curves and regional residuals; when evidence is weak, label as surrogate–target, widen/refresh the oracle slice, and target labels where error concentrates.
@@
AutoCal-R can fail in three distinct ways, each with observable signatures:
@@
The policy-wise mean transport test (3) is a necessary but not sufficient condition for full surrogate transportability (S2): it guarantees that the surrogate is unbiased for policy-wise mean values, but does not rule out conditional miscalibration within subgroups or at the tails of the score distribution. Mitigations: perform subgroup-specific transport tests when fairness concerns apply; monitor reliability diagrams for regional deviations; use richer moment conditions for stronger tests (see remark after 3).
@@
When $\mathbb{E}[Y\mid X,A,S]\neq\mu(S)$, the residual $\Delta(X,A,S)$ induces bias proportional to calibration error. Mitigations: by Cauchy–Schwarz, $|\mathrm{Bias}|\leq\|m-W_{\pi^{\prime}}\|_{2}\,\|\Delta\|_{2}$; DbP shrinks $\|m-W_{\pi^{\prime}}\|_{2}$, so bias is second order when either calibration is tight or the violation small; we surface this via diagnostics and invoke Refuse-Level when unbounded.
@@
OUA assumes the oracle slice is i.i.d. with non-overlapping folds. Mitigations: reuse deterministic folds across modules, de–duplicate the slice, and periodically refresh it.
@@
Scanning many $\pi^{\prime}$ inflates winner’s curse. Mitigations: use FDR control (BH/BY), optionally an outer split for IF–Stack to reduce selection optimism, and emphasize pre–specified contrasts.
@@
Accurate propensities require deterministic, chat–native TF (stable tokenizer/template) with additivity/conditionality invariants; missing/invalid TF corrupts ratios. In our experiments, 39 of 5,000 prompts (0.8%) were filtered: 11 due to base-policy TF conformance failures, and 28 due to missing target-policy logprobs. While small, we report this explicitly: practitioners should expect some data loss from TF gaps, and filtering should always be logged for reproducibility. This also highlights the current state of TF support from LLM providers: major closed-source providers (OpenAI, Anthropic, Google) do not currently support teacher forcing for chat models, and among open-weight hosting providers, Fireworks AI was the only service we found with reliable TF support. Even for our TF reliability test (clone = base with different seed), raw ESS is 26% rather than the theoretical 100%, reflecting API-level non-determinism and floating-point divergence in log-probability summation; SIMCal-W restores this to 99% (Table 4). This 26% is still substantially higher than other policies (0.4–0.7%), confirming TF captures real distributional similarity. The dominant issue is CLE, not TF quality: LLM trajectories are high-dimensional enough that even moderate policy differences yield poor coverage (see Table 4). Mitigations: enforce schema/conformance checks, ledger failures, and treat results as conditional on TF quality (App. E).
@@
DR-CPO adds one rollout + judge per $(X,\pi^{\prime})$; AutoCal–R refits for OUA add modest overhead. We amortize via TF caches, shared folds, and a small stacking library.
@@
Reproducibility Statement. Exact prompts, schema, TF contract, pseudocode, and numerics appear in the appendices. Code and data are available at https://github.com/cimo-labs/cje-arena-experiments; the CJE library is at https://github.com/cimo-labs/cje.
@@ 7 Conclusion
We introduced CJE, a framework that makes it affordable to aim at the right target. By calibrating cheap judges against a small oracle slice, teams no longer need to choose between expensive labels and proxies that may point the wrong direction. CJE fixes all three failures, built around a single rule: Design-by-Projection (DbP). The principle is simple: encode justified assumptions as closed convex sets and project valid objects onto them. Projections onto subspaces (nuisance-orthogonal scores), monotone cones (mean-preserving reward/weight calibration), and simplices (variance-hedged stacking) preserve the estimand while weakly reducing variance.
Concretely, AutoCal-R learns a mean-preserving surrogate $R=f(Z(S))$; SIMCal-W stacks $S$-monotone candidates to produce unit-mean ratios with empirical $\mathrm{ESS}$ gains via OOF stacking and a variance guard; DR-CPO delivers $\sqrt{n}$ inference under cross-fitting; IF-Stack minimizes plug-in IF variance; and OUA with bootstrap and bias-corrected estimation ($\hat{\theta}_{\mathrm{aug}}$) achieves near-nominal CI coverage.
Theoretically, the Efficiency via Model Restriction result, applying established semiparametric principles to surrogate-based evaluation, explains why encoding justified knowledge as model restrictions lowers the efficiency bound, which is attainable with projection-designed estimators and cross-fitting. Two design corollaries follow: (i) Blackwell–efficiency monotonicity: richer judges (finer $\sigma$-fields) strictly help; (ii) under the exact $\mathrm{IsoMeanOne}_{S}$ projection, isotonic mean-one calibration weakly reduces weight dispersion (the reference implementation uses a simpler normalization that empirically yields similar $\mathrm{ESS}$ gains). Empirically, on Arena-derived logs, SIMCal-W turns near-degenerate ratios into stable weights (large $\mathrm{ESS}$ gains), stacked-DR achieves near-nominal coverage and near-$\sqrt{n}$ scaling, and stacking improves ordering; when calibration support is limited, CJE flags the issue and reports robust rankings with conservative uncertainty (Refuse-Level).
@@
(i) Default to Direct: OPE fails under low overlap even with high ESS; calibrated judge scores suffice when transportability holds.
@@
Selection-aware inference over large policy sets; robust/DP isotonic calibration (mirror/Bregman DbP) for heavy tails; active oracle budgeting via shadow prices; sequential/agent evaluations with prefix-aware SIMCal and stepwise DR; and subgroup-aware constraints with fairness diagnostics.
@@ 8 Related Work
Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002; Niculescu-Mizil & Caruana, 2005), and temperature scaling (Guo et al., 2017) are standard post-hoc calibration methods for classification. Our AutoCal-R applies isotonic regression not for classification confidence but for reward calibration: mapping a continuous judge score $S$ to an expected oracle outcome $\mathbb{E}[Y\mid S]$. The mean-preserving property (Lemma 1) ensures the expected value of calibrated rewards matches the oracle under the logging distribution.
@@
IPS (Horvitz & Thompson, 1952; Hájek, 1964) and doubly robust estimators (Bang & Robins, 2005; Jiang & Li, 2016; Kallus & Uehara, 2020) are foundational. Weight stabilization via clipping (Ionides, 2008), balancing (Kallus, 2018), and overlap weighting (Li et al., 2018; Crump et al., 2009) address variance inflation. Most relevant is recent work on isotonic-calibrated IPW (van der Laan et al., 2025) and DR inference via calibration (van der Laan et al., 2024). Our SIMCal-W extends this by projecting importance weights onto the cone of judge-score-monotone functions, exploiting surrogate structure absent in generic OPE.
@@
Super Learner (van der Laan et al., 2007) provides a powerful framework for combining candidate algorithms by minimizing cross-validated prediction risk. Recent work extends ensemble ideas to semiparametric estimation (Shin et al., 2020), combining efficient estimators derived from factored likelihoods. Our IF-space stacking (Theorem 5) takes a complementary approach: instead of combining predictions to minimize MSE, we combine estimators by minimizing the asymptotic variance of their combined influence function. This provides a general formulation that directly targets efficiency, is not tied to specific likelihood factorizations, and is explicitly framed as optimization in the Hilbert space of influence functions.
@@
Generalizing from source to target populations is formalized by the selection diagram framework (Bareinboim & Pearl, 2013a; b; Pearl & Bareinboim, 2014), which uses selection variables $\mathbf{S}$ to encode mechanisms that differ between domains. A quantity is transportable when selection nodes can be “blocked” by conditioning (graphically, when $\mathbf{S}\perp\!\!\perp Y\mid(X,Z)$ in the manipulated graph). Our S2 assumption instantiates this: calibration transports when no selection mechanism points into $Y$ given $(X,A,S)$. The dashed red edges in Figure 1 represent potential transport failures: policy shift ($\pi\to Y$) or environment shift ($E\to Y$) can break the $S$-$Y$ calibration, and the transport test detects when this occurs. Covariate shift methods (Shimodaira, 2000; Sugiyama et al., 2007) address the special case where selection affects only covariates ($E\to X$), not outcomes directly. The CLE bound (Section 2) formalizes a complementary failure mode: even with valid transportability, logs-only estimation fails under poor coverage, motivating Direct or DR methods.
@@ Acknowledgements
We thank Kevin Zielnicki, Molly Davies, Sven Schmit, Brad Klingenberg, Chris Rinaldi, Viridiana Lourdes, Izzy Farley, Sudhish Kasaba Ramesh, Sean Kelly, Adith Swaminathan, Jasmine Nettiksimmons, Winston Chou, Manjari Narayan, and Tom Cunningham for helpful discussions and feedback on earlier drafts of this work.
@@ A.1 Assumptions (compact)
(J2-M) Mean sufficiency (monotone). $\mathbb{E}[Y\mid X,A,S]=\mu(S)$ with $\mu$ weakly nondecreasing. (J2-SI) Single-index fallback. There exist $g^{\star}:\mathcal{S}\times\mathcal{X}\!\to\!\mathbb{R}$ and nondecreasing $\mu^{\star}$ such that $\mathbb{E}[Y\mid X,A,S]=\mu^{\star}\!\big(g^{\star}(S,X)\big)$. When $g^{\star}$ depends only on $S$, this reduces to J2-M; when covariates $X$ (e.g., response length) are included in $g$, the target remains $V(\pi^{\prime})$ and the estimand is preserved provided the index is consistently estimated.
@@
(R2) Calibration consistency. AutoCal-R satisfies $\|\hat{f}(Z)-\mathbb{E}[Y\mid Z]\|_{L^{2}(P_{Z})}=o_{p}(1)$, where $Z=S$ in monotone mode and $Z=g(S,X)$ (learned index) in two-stage mode. SIMCal-W satisfies $\|\hat{W}_{\pi^{\prime}}-m^{\star}_{\rho}\|_{L^{2}(P)}=o_{p}(1)$, where $m^{\star}_{\rho}$ is the $S$-monotone, mean-one, variance-capped projection of $W_{\pi^{\prime}}$ (see Section A.3); when $\mathbb{E}[W_{\pi^{\prime}}\mid S]$ is itself monotone and satisfies the variance constraint, $m^{\star}_{\rho}=\mathbb{E}[W_{\pi^{\prime}}\mid S]$.

### v3

@@ 6 Limitations
We assume the operational oracle $Y$ aligns with stakeholder values; oracle selection itself is a governance question outside the scope of this work (Assumption A0 in App. C.15).
@@
IPS requires support overlap between $\pi_{0}$ and each $\pi^{\prime}$; DR benefits from overlap but remains consistent via its outcome model when overlap is poor; Direct requires no overlap (it generates fresh responses under $\pi^{\prime}$). When overlap is poor, raw ratios are heavy-tailed and uncertainty inflates. Mitigations: Weight stabilization reduces dispersion and raises $\mathrm{ESS}$; if tails persist we (i) gate on $\mathrm{ESS}$ and Hill indices, (ii) use overlap weighting or cohort restriction, and (iii) run an online check when $\widehat{\alpha}_{\text{Hill}}{<}1$ or single–row dominance persists (App. D).
@@
Our calibration model uses a monotone single-index restriction (J2-M/J2-SI) as a stability/regularization choice. Correctness of level claims across policies is governed by mean transport ($\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$), which we audit; when the audit fails or evidence is weak, we widen/refresh the oracle slice or refuse level claims. Mitigations: (i) With audit: run the mean transport test per policy (4); fail $\to$ recalibrate or refuse levels. (ii) Without audit: surface reliability curves as indirect evidence; when evidence is weak, target labels where error concentrates.
@@
Reward calibration can fail in three distinct ways, each with observable signatures:
@@
The policy-wise mean transport test (4) is a necessary but not sufficient condition for full surrogate transportability (S2): it guarantees that the surrogate is unbiased for policy-wise mean values, but does not rule out conditional miscalibration within subgroups or at the tails of the score distribution. Mitigations: perform subgroup-specific transport tests when fairness concerns apply; monitor reliability diagrams for regional deviations; use richer moment conditions for stronger tests (see remark after 4).
@@
When $\mathbb{E}[Y\mid X,A,S]\neq\mu(S)$, the residual $\Delta(X,A,S)$ induces bias proportional to calibration error. Mitigations: by Cauchy–Schwarz, $|\mathrm{Bias}|\leq\|m-W_{\pi^{\prime}}\|_{2}\,\|\Delta\|_{2}$; weight stabilization shrinks $\|m-W_{\pi^{\prime}}\|_{2}$, so bias is second order when either calibration is tight or the violation small; we surface this via diagnostics and invoke Refuse-Level when unbounded.
@@
Calibration-aware inference assumes the oracle slice is i.i.d. with non-overlapping folds. Mitigations: reuse deterministic folds across modules, de–duplicate the slice, and periodically refresh it.
@@
Scanning many $\pi^{\prime}$ inflates winner’s curse. Mitigations: use FDR control (BH/BY), optionally an outer split for influence-function stacking to reduce selection optimism, and emphasize pre–specified contrasts.
@@
Accurate propensities require deterministic, chat–native TF (stable tokenizer/template) with additivity/conditionality invariants; missing/invalid TF corrupts ratios. In our experiments, 39 of 5,000 prompts (0.8%) were filtered: 11 due to base-policy TF conformance failures, and 28 due to missing target-policy logprobs. While small, we report this explicitly: practitioners should expect some data loss from TF gaps, and filtering should always be logged for reproducibility. This also highlights the current state of TF support from LLM providers: major closed-source providers (OpenAI, Anthropic, Google) do not currently support teacher forcing for chat models, and among open-weight hosting providers, Fireworks AI was the only service we found with reliable TF support. Even for our TF reliability test (clone = base with different seed), raw ESS is 26% rather than the theoretical 100%, reflecting API-level non-determinism and floating-point divergence in log-probability summation; weight stabilization restores this to 99% (Table 4). This 26% is substantially higher than other policies (0.4–0.7%), confirming TF captures real distributional similarity, but it is far from the theoretical 100% and suggests propensity estimates carry meaningful error even before overlap considerations. OPE failure in our experiments reflects two compounding issues: (i) TF brittleness corrupts propensity estimates, and (ii) CLE limits precision even with perfect propensities. We cannot cleanly separate their contributions, but both matter: if TF were the only problem, weight stabilization should improve ranking more than it does; if CLE were the only problem, clone should have ESS $\approx$ 100%. Mitigations: enforce schema/conformance checks, ledger failures, and treat results as conditional on TF quality (App. E).
@@
Calibrated DR adds one rollout + judge per $(X,\pi^{\prime})$; reward calibration refits for calibration-aware inference add modest overhead. We amortize via TF caches, shared folds, and a small stacking library.
@@
Reproducibility Statement. Exact prompts, schema, TF contract, pseudocode, and numerics appear in the appendices. Code and data are available at https://github.com/cimo-labs/cje-arena-experiments; the CJE library is available via pip install cje-eval (https://github.com/cimo-labs/cje).
@@ 7 Conclusion
We introduced CJE, a framework that makes it affordable to aim at the right target. By calibrating cheap judges against a small oracle slice, teams no longer need to choose between expensive labels and proxies that may point the wrong direction. CJE fixes all three failures, built around a single principle: encode justified assumptions as model restrictions. Restrictions to subspaces (nuisance-orthogonal scores), monotone cones (mean-preserving reward/weight calibration), and simplices (variance-hedged stacking) preserve the estimand while weakly reducing variance.
Concretely, reward calibration learns a mean-preserving surrogate $R=f(Z(S))$; weight stabilization stacks $S$-monotone candidates to produce unit-mean ratios with empirical $\mathrm{ESS}$ gains via OOF stacking and a variance guard; calibrated DR delivers $\sqrt{n}$ inference under cross-fitting; influence-function stacking minimizes plug-in IF variance; and calibration-aware inference with bootstrap and bias-corrected estimation ($\hat{\theta}_{\mathrm{aug}}$) achieves near-nominal CI coverage.
Theoretically, the Efficiency via Model Restriction result, applying established semiparametric principles to surrogate-based evaluation, explains why encoding justified knowledge as model restrictions lowers the efficiency bound, which is attainable with projection-designed estimators and cross-fitting. Two design corollaries follow: (i) Blackwell–efficiency monotonicity: richer judges (finer $\sigma$-fields) strictly help; (ii) under the exact $\mathrm{IsoMeanOne}_{S}$ projection, isotonic mean-one calibration weakly reduces weight dispersion (the reference implementation uses a simpler normalization that empirically yields similar $\mathrm{ESS}$ gains). Empirically, on Arena-derived logs, weight stabilization turns near-degenerate ratios into stable weights (large $\mathrm{ESS}$ gains), stacked-DR achieves near-nominal coverage and near-$\sqrt{n}$ scaling, and stacking improves ordering; when calibration support is limited, CJE flags the issue and reports robust rankings with conservative uncertainty (Refuse-Level).
@@
(i) Default to Direct: For open-ended generation under nontrivial policy shifts, logs-only OPE faces two compounding issues—TF brittleness and coverage sparsity (low TTC)—that together create a precision floor weight stabilization cannot overcome. In constrained settings or near-clone regimes, OPE may remain viable; gate on TTC $\geq 0.70$ before relying on logged data.
@@
Selection-aware inference over large policy sets; robust/DP isotonic calibration (mirror/Bregman projections) for heavy tails; active oracle budgeting via shadow prices; sequential/agent evaluations with prefix-aware weight stabilization and stepwise DR; and subgroup-aware constraints with fairness diagnostics.
@@ 8 Related Work
Auditable assumptions via negative controls and proxies.
CJE’s approach to transportability draws on a broader principle in causal inference: turning untestable identification conditions into checkable diagnostics using auxiliary data. Negative controls provide a canonical example: variables known to have no causal relationship with treatment or outcome can detect residual confounding when associations appear (Lipsitch et al., 2010). Proximal causal inference extends this idea, using proxy variables to construct testable restrictions that support identification under weaker assumptions (Tchetgen Tchetgen et al., 2024). CJE applies a similar philosophy: rather than assuming global surrogacy, we budget oracle labels in each target context to test a policy-wise moment restriction ($\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$). The contribution is operational: we provide an evaluation protocol with explicit pass/fail criteria and decision rules (reuse calibration vs. recalibrate vs. refuse level claims), making the assumption auditable rather than hoping it holds. CJE is not a negative-control estimator; the parallel is conceptual and design-oriented, not methodological.
@@
Misclassification correction.
Rogan & Gladen (1978) introduced prevalence estimation under measurement error; Lee et al. (2025) adapt this for LLM judges, providing confidence intervals that account for calibration uncertainty. Their approach assumes confusion-matrix stability—that sensitivity and specificity transfer from calibration to test—analogous to our transport condition. We compare directly in Section 5.3 and find continuous calibration $13\times$ more sample-efficient, even when binary correction is numerically well-conditioned (mean $J=0.41$). The efficiency gap arises from ill-conditioning (the Rogan-Gladen estimator’s standard error scales like $1/J$) and information loss from binary thresholding.
EIF-based debiasing.
Concurrent work by Chen et al. (2026) formalizes LLM-as-judge debiasing through efficient influence functions (EIF), unifying Rogan–Gladen-style measurement error correction with PPI/PPI++-style calibration. They derive the EIF for mean estimation with a surrogate and labeled subset, showing that naive judge-based CIs can achieve 0% coverage under bias while EIF-based estimators maintain nominal coverage with substantially narrower intervals than confusion-matrix methods. Their framing—learning $\mu(\hat{Y})=\mathbb{E}[Y\mid\hat{Y}$] on a small calibration set and using residual augmentation—is closely aligned with CJE’s reward calibration. CJE extends this to continuous scores, policy/deployment shift (via the mean transport test), and off-policy evaluation with weight stabilization.
@@
Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002; Niculescu-Mizil & Caruana, 2005), and temperature scaling (Guo et al., 2017) are standard post-hoc calibration methods for classification. Our reward calibration applies isotonic regression not for classification confidence but for reward calibration: mapping a continuous judge score $S$ to an expected oracle outcome $\mathbb{E}[Y\mid S]$. The mean-preserving property (Lemma 1) ensures the expected value of calibrated rewards matches the oracle under the logging distribution.
@@
IPS (Horvitz & Thompson, 1952; Hájek, 1964) and doubly robust estimators (Bang & Robins, 2005; Jiang & Li, 2016; Kallus & Uehara, 2020) are foundational. Weight stabilization via clipping (Ionides, 2008), balancing (Kallus, 2018), and overlap weighting (Li et al., 2018; Crump et al., 2009) address variance inflation. Most relevant is recent work on isotonic-calibrated IPW (van der Laan et al., 2025) and DR inference via calibration (van der Laan et al., 2024). Our weight stabilization extends this by projecting importance weights onto the cone of judge-score-monotone functions, exploiting surrogate structure absent in generic OPE.
@@
Super Learner (van der Laan et al., 2007) provides a powerful framework for combining candidate algorithms by minimizing cross-validated prediction risk. Recent work extends ensemble ideas to semiparametric estimation (Shin et al., 2020), combining efficient estimators derived from factored likelihoods. Our influence-function stacking (Theorem 5) takes a complementary approach: instead of combining predictions to minimize MSE, we combine estimators by minimizing the asymptotic variance of their combined influence function. This provides a general formulation that directly targets efficiency, is not tied to specific likelihood factorizations, and is explicitly framed as optimization in the Hilbert space of influence functions.
@@
Generalizing from source to target populations is formalized by the selection diagram framework (Bareinboim & Pearl, 2013a; b; Pearl & Bareinboim, 2014), which uses selection variables $\mathbf{S}$ to encode mechanisms that differ between domains. A quantity is transportable when selection nodes can be “blocked” by conditioning (graphically, when $\mathbf{S}\perp\!\!\perp Y\mid(X,Z)$ in the manipulated graph). Our transport assumption instantiates this: calibration transports when no selection mechanism points into $Y$ given $(X,A,S)$. The dashed red edges in Figure 1 represent potential transport failures: policy shift ($\pi\to Y$) or environment shift ($E\to Y$) can break the $S$-$Y$ calibration, and the transport test detects when this occurs. Covariate shift methods (Shimodaira, 2000; Sugiyama et al., 2007) address the special case where selection affects only covariates ($E\to X$), not outcomes directly. The CLE bound (Section 2) formalizes a complementary failure mode: even with valid transportability, logs-only estimation fails under poor coverage, motivating Direct or DR methods.
@@ Acknowledgements
We thank Kevin Zielnicki, Molly Davies, Sven Schmit, Brad Klingenberg, Chris Rinaldi, Viridiana Lourdes, Izzy Farley, Sudhish Kasaba Ramesh, Sean Kelly, Adith Swaminathan, Jasmine Nettiksimmons, Winston Chou, and Tom Cunningham for helpful discussions and feedback on earlier drafts of this work.
@@ A.1 Assumptions (compact)
(J2-M) Mean sufficiency (monotone). $\mathbb{E}[Y\mid X,A,S]=\mu(S)$ with $\mu$ weakly nondecreasing. This structural assumption enables the EIF derivations (Theorem 1); it is sufficient for transport (T1) but not required when transport is verified via audit. (J2-SI) Single-index fallback. There exist $g^{\star}:\mathcal{S}\times\mathcal{X}\!\to\!\mathbb{R}$ and nondecreasing $\mu^{\star}$ such that $\mathbb{E}[Y\mid X,A,S]=\mu^{\star}\!\big(g^{\star}(S,X)\big)$. When $g^{\star}$ depends only on $S$, this reduces to J2-M; when covariates $X$ (e.g., response length) are included in $g$, the target remains $V(\pi^{\prime})$ and the estimand is preserved provided the index is consistently estimated.
(T1) Transport. $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$. Calibration learned under $\pi_{0}$ is mean-unbiased under $\pi^{\prime}$. This is the fundamental requirement for valid surrogate-only policy value estimation. With an audit slice under $\pi^{\prime}$, T1 is testable (4); without audit, T1 must be assumed (e.g., via J2-M/J2-SI or domain knowledge). See the assumption hierarchy in Section C.15.
@@
(R2) Calibration consistency. Reward calibration satisfies $\|\hat{f}(Z)-\mathbb{E}[Y\mid Z]\|_{L^{2}(P_{Z})}=o_{p}(1)$, where $Z=S$ in monotone mode and $Z=g(S,X)$ (learned index) in two-stage mode. Weight stabilization satisfies $\|\hat{W}_{\pi^{\prime}}-m^{\star}_{\rho}\|_{L^{2}(P)}=o_{p}(1)$, where $m^{\star}_{\rho}$ is the $S$-monotone, mean-one, variance-capped projection of $W_{\pi^{\prime}}$ (see Section A.3); when $\mathbb{E}[W_{\pi^{\prime}}\mid S]$ is itself monotone and satisfies the variance constraint, $m^{\star}_{\rho}=\mathbb{E}[W_{\pi^{\prime}}\mid S]$.